OpenStudy (hdrager):

please help I totaly forgot how to do this solve xy+y+1=y for y

Nnesha (nnesha):

get all y terms on the right side...

OpenStudy (hdrager):

i know i just forgot how to do that without making a mess of the equation

Nnesha (nnesha):

left*

OpenStudy (kittiwitti1):

\[xy+y+1=y\]\[xy+y\cancel{+1-1}-y=\cancel{y-y}-1\]\[xy+y-y=-1\]

OpenStudy (kittiwitti1):

can you do it from here?

OpenStudy (hdrager):

y=-1/x

OpenStudy (hdrager):

is that it?

OpenStudy (kittiwitti1):

yes

OpenStudy (kittiwitti1):

is this multiple choice or free response?

OpenStudy (hdrager):

it doesn't seem complete to me...

OpenStudy (hdrager):

it's free response

OpenStudy (hdrager):

okay cool thanks

OpenStudy (kittiwitti1):

well if they give you two variables and ask for one, then it will end up being in terms of the other variable

OpenStudy (kittiwitti1):

but if you're not sure you can ask someone else :p

OpenStudy (hdrager):

i just have a bad habit of starting a problem and thinking i've finished when i haven't

OpenStudy (hdrager):

so i try to be super cautious of it

OpenStudy (kittiwitti1):

ah I see